# Aleph Filter: To Infinity in Constant Time

### v1

@@ 1. Introduction
term
definition
$N$
current filter capacity divided by initial capacity
$X$
number of expansions so far (i.e., $X=\lceil\log_{2}(N)\rceil$)
$M$
total memory used for the filter (bits / entry)
$F$
initial fingerprint length when the filter is first allocated
$h(...)$
mother hash generating function
$\alpha$
fraction of occupied slots ($10\leq\alpha<$ 1)
@@ 2.2. InfiniFilter
query/ delete/rejuv
insert
FPR
fingerprint
@@
Fingerprint Sacrifice (Zhang et al., 2021; Bender et al., 2012)
$O(1)$
$O(1)$
$O(2^{-F}\cdot N)$
$F-O(\lg N)$
$F$
InfiniFilter (Fixed-Width Regime)
$O(\frac{\lg N}{F})$
$O(1)$
$O(2^{-F}\cdot\lg N)$
$F$
$2^{F}$
InfiniFilter (Widening Regime)
$O(\frac{\lg N}{F+\lg\lg N})$
$O(1)$
$O(2^{-F})$
$F+O(\lg\lg N)$
$\infty$
@@ 4.1. Fast Queries by Duplicating Void Entries
query/ delete/rejuv
insert
false positive rate
fingerprint
@@
Aleph Filter (Fixed-Width Regime)
$O(1)$
$O(1)$
$O(2^{-F}\cdot\lg N)$
$F$
$2^{F}$
Aleph Filter (Widening Regime)
$O(1)$
$O(1)$
$O(2^{-F})$
$F+O(\lg\lg N)$
$\infty$
Aleph Filter (Predictive Regime)
$O(1)$
$O(1)$
$O(2^{-F})$
$F+O(\lg|\lg(N/N_{est})|)$
$\infty$
@@ 6. Related Work
Theoretical Algorithms and Bounds. Pagh, Segev, and Wieder prove a lower bound that if we initialize a filter to constant capacity and expand it to contain $N$ keys, the filter must at some point use at least $\lg\lg N$ bits per key in addition to what is required for a filter with a fixed capacity of $N$ keys (Pagh et al., 2013). Aleph Filter’s Widening and Predictive Regimes both meet this lower bound. The former meets it after the $N$ data entries are inserted, while the latter meets it before any entries are inserted.
The idea of duplicating void entries within a filter hash table to support indefinite expansion with constant time queries can also be traced back to Pagh, Segev, and Wieder (Pagh et al., 2013). Nevertheless, their algorithm has never been implemented, and Aleph filter improves on it in four important ways. (1) Aleph filter requires $O(\lg\lg N)$ rather than $O(\lg\lg U)$ bits per key to support indefinite growth while maintaining a stable false positive rate. This is a significant difference if $U$, the universe size, is unknown or significantly larger than $N$, the data size. (2) The data structure in (Pagh et al., 2013) keeps track of the age entries by encoding a binary age counter alongside each slot for $\lg\lg N$ additional bits per key. In contrast, Aleph filter keeps track of the ages of entries using a chain of exponentially smaller hash tables that take up negligible memory. (3) The data structure in (Pagh et al., 2013) handles fingerprint collisions by storing the full keys of entries with colliding fingerprints in an auxiliary dictionary. In contrast, Aleph filter supports storing multiple identical fingerprints in the main hash table by employing an underlying quotient filter with Robin Hood hashing. This eliminates one dictionary access from the query path. (4) Aleph Filter introduces the Predictive Regime, which can drastically reduce the memory requirement.

### v2

@@ 1. Introduction
term definition
$N$ current filter capacity divided by initial capacity
$X$ number of expansions so far (i.e., $X=\lceil\log_{2}(N)\rceil$)
$M$ total memory used for the filter (bits / entry)
$F$ initial fingerprint length when the filter is first allocated
$h(...)$ mother hash generating function
$\alpha$ fraction of occupied slots ($10\leq\alpha<$ 1)
@@ 2.2. InfiniFilter
query/ delete/rejuv insert FPR fingerprint
@@
Fingerprint Sacrifice (Zhang et al., 2021; Bender et al., 2012) $O(1)$ $O(1)$ $O(2^{-F}\cdot N)$ $F-O(\lg N)$ $F$
InfiniFilter (Fixed-Width Regime) $O(\frac{\lg N}{F})$ $O(1)$ $O(2^{-F}\cdot\lg N)$ $F$ $2^{F}$
InfiniFilter (Widening Regime) $O(\frac{\lg N}{F+\lg\lg N})$ $O(1)$ $O(2^{-F})$ $F+O(\lg\lg N)$infty
@@ 4.1. Fast Queries by Duplicating Void Entries
query/ delete/rejuv insert false positive rate fingerprint
@@
Aleph Filter (Fixed-Width Regime) $O(1)$ $O(1)$ $O(2^{-F}\cdot\lg N)$ $F$ $2^{F}$
Aleph Filter (Widening Regime) $O(1)$ $O(1)$ $O(2^{-F})$ $F+O(\lg\lg N)$infty
Aleph Filter (Predictive Regime) $O(1)$ $O(1)$ $O(2^{-F})$ $F+O(\lg|\lg(N/N_{est})|)$infty
@@ 6. Related Work
From Theory into Practice. Pagh, Segev, and Wieder prove a lower bound that if we initialize a filter to constant capacity and expand it to contain $N$ keys, the filter must at some point use at least $\lg\lg N$ bits per key in addition to what is required for a filter with a fixed capacity of $N$ keys (Pagh et al., 2013). Aleph Filter’s Widening and Predictive Regimes both meet this lower bound. The former meets it after the $N$ data entries are inserted, while the latter meets it before any entries are inserted.
The idea of duplicating void entries within a filter hash table to support indefinite expansion with constant time queries can also be traced back to Pagh, Segev, and Wieder (Pagh et al., 2013). Taffy Cuckoo Filter (Apple, 2022) is an expandable Cuckoo filter based on these ideas, though it does not support deletes or rejuvenation operations.
Aleph filter improves on these works in several ways. (1) It requires $O(\lg\lg N)$ rather than $O(\lg\lg U)$ bits per key to support indefinite growth while maintaining a stable false positive rate. This is a significant difference if $U$, the universe size, is unknown or significantly larger than $N$, the data size. (2) To support deletes, the data structure in (Pagh et al., 2013) keeps track of the age of entries by encoding a binary age counter alongside each slot for $\lg\lg N$ additional bits per key. In contrast, Aleph filter keeps track of the ages of entries using a chain of exponentially smaller hash tables that take up negligible space. (3) The data structure in (Pagh et al., 2013) handles fingerprint collisions by storing the full keys of entries with colliding fingerprints in an auxiliary dictionary. In contrast, Aleph filter supports storing multiple identical fingerprints in the main hash table by employing an underlying quotient filter with Robin Hood hashing. This eliminates one dictionary access from the query path. (4) Aleph Filter introduces the Predictive Regime, which can drastically reduce the memory requirement.